RC Meeting and Runplan, Friday, Sept 21
- Target is setup so that selection of liquid targets is disabled.
- Current program is establishing beam to Halls C and D, simultaneously. Beam is highly unlikely in our hall until just after midnight. Slim possibility at the end of Swing shift tonight. We will begin with Jay's beam deflection test. SHMS and HMS are set at the first angles we need (6.66 and 13.5 deg, respectively). SHMS was ramped today to -8.035 GeV and back to 0 for testing of procedures and magnets.
- To clarify the run plan going forward and this evening:
1. Ion chamber calibration, BPM calibration, the usual ops stuff etc... 2. Ops is trying to establish high current beam to C. However, they will likely stop at some point and send us what they can. We can do trigger and DAQ commissioning if that's the case. 3. The beam deflection test is still the first item on our agenda, but it is not a show stopper, and we can continue forward and go back to it if we encounter a problem or have a weird result. The beam deflection test is linked here: